OpenStudy (anonymous):

using change of variable: lim h->0 square root of x -2 / square root of x cubed -2

OpenStudy (anonymous):

\[\sqrt{x}-2 / \sqrt{x ^{3}} -8\]

OpenStudy (anonymous):

sorry wrote the question wrong.

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0} (\sqrt{(x-2)/\sqrt{(x^3 -2)}})\]

OpenStudy (anonymous):

the question is \[\sqrt{x}-2/\sqrt{x ^{3}} -8\]

OpenStudy (anonymous):

+/- (1/2)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}[\sqrt{x-2}/\sqrt{x-2}*\sqrt{x^2+2x+4}]\] \[\lim_{x \rightarrow 0}[1/\sqrt{x^2+2x+4}]\] when you plug in the value of x=0 \[= 1/\sqrt{4}\] \[+/- (1/2)\]